OpenStudy (anonymous):

Graph the two equations to solve the system. What is the solution? 3x - 2y = 6 and 5x + 3y = -9 A. (4, 3) B. (5, 4.5) C. (0, -3) D. infinitely many solutions

OpenStudy (mertsj):

C. Plug the points in and see which one makes both sentences true.

OpenStudy (radar):

I would plot the two lines by locating their two respective intercepts by letting x=0 then y=0

OpenStudy (mertsj):

@Luis They don't even have the same slope. How could they be the same line and have infinitely many solutions?

OpenStudy (anonymous):

ty

OpenStudy (mertsj):

yw

OpenStudy (radar):

3x-2y=6 when x=0, y=-3 (0,-3) when y=0 x=2(2,0) 5x+3y=-9, when x=0, y=-3 (0,-3) when y=0, x=-9/5) Now draw each line using the two points for each respective line. If and where the lines intersect, that will be the answer

OpenStudy (mertsj):

@ Luis Well, they are not parallel but even if they were, there would be NO solution not infinitely many.

OpenStudy (mertsj):

@ Luis The slope of the first one is 3/2. The slope of the second is -5/3 so they are clearly not parallel.

OpenStudy (radar):